OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

perpendicular means inverse opposite slope

OpenStudy (anonymous):

So, y = -1/2x + -2?

OpenStudy (anonymous):

Oops I meant y = -1/2x -2

OpenStudy (anonymous):

no you need to use the points 6,3

OpenStudy (anonymous):

y-3=-1/2(x-6)

OpenStudy (anonymous):

I think I get it.

OpenStudy (anonymous):

What would the equation look like out of these? y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3

OpenStudy (shamil98):

you could just plug in the values and solve for b. y = -1/2x + b 3 = -1/2(6)+ b 3 = -3 + b 6 = b

OpenStudy (anonymous):

OOOHH So, then It'd be the third answer?

OpenStudy (shamil98):

yes